Maximize with respect to p. What value of p do you get?

Maximize

with respect to p. What value of p do you get?
Answer #1

Let f( ) = log (CP3(1-p)2) = log ( df(p) 32 t. 3 log p +2 log(1-P) dp P 1-p p -=-p2 _ (i-p)2 < 0 for p = 5 Hence at p =-, f(p